OpenStudy (anonymous):

I forget how to do these. (3 square root 2)(2 square root 6)=?

OpenStudy (anonymous):

\[\sqrt{a} \times \sqrt{b} = \sqrt{ab}\]

OpenStudy (anonymous):

i see! thx